If the sides of a square are increased by 2 meters, the area becomes 49 square meters. Find the length of a side of the original square.
Answer by ankor@dixie-net.com(22740)

If the sides of a square are increased by 2 meters, the area becomes 49 square meters.
Find the length of a side of the original square.
;
Let x = side of the original square
then
(x+2) = side of the larger square
:
Find the area of the larger square
(x+2)*(x+2) = 49
FOIL
x^2 + 4x +4 = 49
:
x^2 + 4x + 4 - 49 = 0
:
x^2 + 4x - 45 = 0; a quadratic equation
Factors to:
(x+9)(x-5) = 0
Positive solution is what we want here
x = 5 meters is the side of the original square
;
;
Check solution: (5+2)^2 = 49